Can you help me solve this problem and show me the steps to find the answer? 50 dogs, 10 are white fur recessive, 40 are blue fur dominant. The homozgous is .70. What is the ratio dogs that are heterozgous?


1
Expert's answer
2021-03-29T13:17:29-0400

According to the Hardy-Weinberg equation:

p2 + 2pq + q2 = 1

If 0.70 dogs are homozygous, the ratio of heterozygous dogs equals:

2pq = 1 - (p2 + q2) = 1 - 0.7 = 0.3


Answer: 0.3